War Game, Inc., produces historical battle simulation games in a small market. It is thinking about introducing a new game in honor of the seventy-fifth anniversary of the outbreak of World War II. Based on historical data, management forecasts demand for this game to be P = 800 - .008Q where Q denotes unit sales per year, and P denotes price in dollars. The cost of manufacture (based on royalty payments to the designer of the game, and the costs of printing and distributing) is C = 120,000 + 10Q

a. If the goal of War Game is to maximize profit, calculate the optimal output and price.

b. If instead the company's goal is to maximize sales revenue, what is its optimal price and quantity?
Answer #1

P: 800 -0.008 0 SO Revenue - PQ » P Q = 800 Q-0.008 Q2 MR = a (TR) = d (PQ) dQ - d (800 -0.0089%) do - 800 -0.0160. C = 120,0b) At maximum total revenue - HP = 0 > 800 -0.01680 - 0 = 800/6.01 - 150,000 units P-800- 0.0088 - 800 - 0.00 8 X 50,000 iduo